Johnnyneeds 10liters of a 15% acid solution. He currently has a 10%solution and a 30%solution. How many liters of 10% solution does he need to make the needed 10liters of 15% acidsolution?
Found 2 solutions by Alan3354, htmentor:
Answer by Alan3354(69443)

Johnny needs 10 liters of a 15% acid solution. He currently has a 10% solution and a 30% solution. How many liters of 10% solution does he need to make the needed 10 liters of 15% acid solution?
------------------------
He currently has --- why not just "he has" ---- Currently is not needed.
==========================
x = amount of 10% to use
y = amount of 30%
----
x + y = 10 ---- total solution
10x + 30y = 10*15 ---- total acid
--------------
Solve for x & y

Answer by htmentor(1343)

Let x = the number of liters of the 10% solution. Then 10-x is the volume of the
30% solution.
The final mixture is 10 liters, of which 15% is acid.
So the final mixture has 0.15*10 = 1.5 liters of acid
1.5 = 0.1x + 0.3(10-x) -> 1.5 = 0.2x -> x = 7.5 liters